OpenStudy (anonymous):

Choose the correct simplification of x^2/x^7. A. x^5 B. x^9 C. 1/x^5 D. x^7*y^12/x^4*y^9

OpenStudy (anonymous):

Expand it out and then cancel.

OpenStudy (anonymous):

\[\dfrac{x \cdot x}{x \cdot x \cdot x \cdot x \cdot x \cdot x \cdot x}\]

OpenStudy (anonymous):

\(\dfrac{x}{x} = 1\) so cancel out 1 x on the bottom for every x on the top.